Tom's Uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The 2nd longest side is 2 yards shorter than the longest side. What is the length of each side?
Answer by checkley77(12844)

378=(30+L/2)+(L-2)+L
378=30+L/2+L-2+L
378=2.5L+28
2.5L=378-28
2.5L=350
L=350/2.5
L=140 FOR THE LONGEST SIDE.
30+140/2=30+70=100 FOR THE NEXT SIDE.
140-2=138 FOR THE THIORD SIDE.
PROOF:
140+100+138=378
378=378